OpenStudy (anonymous):

2√(x+1) - [6/√(2x-1)] =0 Find x.

OpenStudy (dmezzullo):

first add \(\frac{6}{\sqrt{2x + 1}}\) to both sides, and then square both. you will be left with no radicals

OpenStudy (anonymous):

what?

OpenStudy (dmezzullo):

can you add \(\frac{6}{\sqrt{2x -1}}\) to both sides?

OpenStudy (anonymous):

2√(x+1) = [6/√(2x-1)]

OpenStudy (dmezzullo):

that's right. now square both sides

OpenStudy (anonymous):

4x+4 = 36/ (2x-1) ?

OpenStudy (anonymous):

x=2 or x= -2.5 (rej.) ??

hartnn (hartnn):

\(\huge \color{Pink}\checkmark \)

OpenStudy (anonymous):

@dmezzullo never mind, you have tried your best:) you have helped me already:) thanks:) @hartnn thanks:)

OpenStudy (anonymous):

2√(x+1) - [6/√(2x-1)] =0 2√(x+1)=[6/√(2x-1)] squaring both sides we get; 4(x+1)=36/(2x-1) cross multiplication you will get; 4(x+1)(2x-1)=36 (4x+4)(2x-1)=36 8x^2-4x+8x-4=36 8x^2+4x-4-36=0 8x^2+4x-40=0 4(2x^2+x-10)=0 2x^2+x-10=0 on solving you will get x=2 and x= (-2.5 ;this value is neglected as x cant be negative) therefore the value of x=2 is your answer (Note:you can check your answer by putting the value of x=2 in your given equation as well)